OpenStudy (anonymous):

Fran Massey earns three times as much as an actuary then she does as a writer. her total income is $40000 more than that of her brother who earns only half as much as Fran does an actuary. What is Frans salary as an actuary?

OpenStudy (anonymous):

lets say for example that her income was \($36,000\) how much did she make as an actuary, and how much did she make as a writer?

OpenStudy (anonymous):

if it is not clear, let me know and i will see if i can explain it

OpenStudy (anonymous):

she made 12,00 as a writer i believe

OpenStudy (anonymous):

not quite i think you divide by 3 but if she made $12,000 as a writer, and made three times as much as an actuary, then she made \(3\times 12000=36000\) as an actuary for a total of \(12,000+36,000=48,000\) so that is not right

OpenStudy (anonymous):

the correct answer is \(9,000\) because if you multiply that by 3 you get \(27,000\) for a total of \(36,000\)

OpenStudy (anonymous):

in other words, this line Fran Massey earns three times as much as an actuary then she does as a writer. can be interpreted this way: if she makes \(x\) amount as a writer, then she makes \(3x\) as an actuary for a total of \(4x\)

OpenStudy (anonymous):

her brother makes half as much as she does as an actuary if she makes \(3x\) as an actuary, then her brother makes \(\frac{3}{2}x\) total

OpenStudy (anonymous):

ok that was wrong, sorry if she make \(40,000\) more than her brother, we can set \[4x=\frac{3}{2}x+40,000\] and solve for \(x\)

OpenStudy (anonymous):

do you know how to solve this? it takes two steps

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

let me know what you get then we can finish the problem

OpenStudy (anonymous):

16000

OpenStudy (anonymous):

k good then that is the amount she makes as a writer, she makes 3 times that much as an actuary which is what you are asked for

OpenStudy (anonymous):

so she makes 40000

OpenStudy (anonymous):

no she makes \(16,000\) as a writer and so \(3\times 16,000\) as an actuary

OpenStudy (anonymous):

i get \($48,000\) as an actuary

OpenStudy (anonymous):

this was kind of a complicated word problem there are probably other ways to solve, but i don't think any are easier

OpenStudy (anonymous):

yeah i got 48000 when i checked it again

OpenStudy (anonymous):

oh good

OpenStudy (anonymous):

alright well thank you for your help